Previous | Calendar |
October |
2006 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
1JD274 (breezy)Site Visit (SpB): Details , Photos (14) added s-h modems to ucb logger, now running "ucb_07.csi" ingest in ASCII format (not binary), start ops30. |
2JD275Yom Kippur |
3JD276 |
4JD277 |
5JD278Site Visit (SpB, JH, MP): Details , Photos (14) |
6JD279Site Visit (SpB): Details , Photos (8) (Full "Harvest" Moon) added s-h modems to tc23x logger, now running "tc_10.csi" ingest tc23x logger data in ASCII format (chan211), start ops31. at west robot: done=2, 4-flasks swapped Summer Jeep Rental (ends, at 7pm) |
7JD280 |
8JD281 |
9JD282 (rain/snow)Columbus Day |
10JD283 |
11JD284 |
12JD285Site Visit (SpB, JH, MP, DM, NT): Details , Photos (1) moved s-h modems to cr21x logger (for pressure sensors) ingest 21x logger data in ASCII format (chan211), start ops32 (note, 21x and 23x ASCII format is different). removed s-h modems from tc23x logger, now running "tc_11.csi" at west robot: done=0, no flasks swapped (fudge=2) Jeep Rental |
13JD286Site Visit (SpB): Details , Photos (16) reconnected the repaired am25t (sn 1396) to tc23x logger (ie, am25t on the ground) modified tc23x logger program for am25t data, now running "tc_12.csi" Jeep Rental |
14JD287Site Visit (SpB, DB, SS, DM, WM, B): Details , Photos (9) at west robot: done=2, swapped 4 flasks (changed fudge=2 to 0) |
15JD288Site Visit (DB, SS): Details |
16JD289Site Visit (SpB, SS): Details , Photos (21) connected more thermocouples to the tc23x am25t set up the "snowpack" ati 2-d sonic (still not on data system) added approx -200mV offset to co2.ec and changed sample range to +/- 1000mV. Ameriflux Meeting |
17JD290 (snow)Ameriflux Meeting |
18JD291Ameriflux Meeting |
19JD292 |
20JD293 |
21JD294 (snow) |
22JD295Site Visit (SpB, LL, MH): Details , Photos (11) AIRCOA is removed from como creek. at west robot: done=1, swapped 4 flasks |
23JD296~21:00 MST, data system crashed (i was messing with minicom!). |
24JD297Site Visit (SpB, JH, MP, DM, NT, RM, DS, S, A): Details , Photos (30) ~10:00 MST, data system back up. at 10:34 MST, increased offset on co2.ec (li-6262) by -100mV. from 10:51-11:04 MST, dessicant swap at li-6251 at 13:49 MST, switched pdiff1 and pdiff2 removed fast ozone stuff from fast logger (fast_02.csi), removed hydra data, and started collecting 2d ati data (chan212), start ops33. Also, renamed a few parameters in prep.config Compact Rental |
25JD298 |
26JD299 (snow)at ~9:15 MST, a brief power outage (several power burps throughout the morning) |
27JD300 |
28JD301 |
29JD302 |
30JD303 |
31JD304Halloween |
1JD305 |
2JD306Site Visit (SpB): Details , Photos (27) connected more thermocouples to the tc23x am25t at west robot: done=1, swapped 4 flasks Compact/Prius Rental |
3JD307 |
4JD308 |